# Chern character of Schur bundles

### v1

@@ 3. Chern character of Schur bundles
$\mathsf{C}^{E}\colon\Lambda\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E).$ (3.1)
It is easy to check that it is a characteristic class as in Definition 1.1. Notice that in this case we have full additivity, not just same-degree additivity (cf. Remark 1.2).
We also point out that
@@
where the last equality is the semi-standard tableaux hook length formula.
Remark 3.1.
By the splitting principle we have $\mathsf{C}_{0}^{E}(\mathsf{\mathbf{p}}_{n})=r$ for every $n\geq 1$. This is the only point where the splitting principle is needed.
@@
However, $\sum_{k}\mathsf{\mathbf{p}}_{k}\frac{\partial}{\partial\mathsf{\mathbf{p}}_{k}}$ acts on homogenous elements via multiplication by their degree (enough to check on a homogeneous basis, like for example the monomials in power sums, where it is evident), thus
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{\alpha}E)=|\alpha|\frac{r_{\alpha}}{r}\operatorname{ch}_{1}(E).$
Remark 3.4.
Corollary 3.3 gives an alternative proof of the main result of [Rub13].
We wish to point out the efficiency of Theorem 3.2 by explicitly computing some examples.
@@
Let us compute $\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E)$ when $|\alpha|=2$. We have only two cases, namely the symmetric power $\mathsf{\mathbf{S}}^{2}E$ and the exterior power $\bigwedge^{2}E$, corresponding to $\alpha=(1,1)$.
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2})=\mathsf{\mathbf{s}}_{1}+2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2})=1.$
@@
$\operatorname{rk}(\mathsf{\mathbf{S}}^{2}E)=$ $\,\frac{r(r+1)}{2}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\operatorname{ch}_{2}(E)+\frac{1}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{1}{2}\operatorname{ch}_{2}(E)^{2}.$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1)})=\mathsf{\mathbf{s}}_{1}-2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1)})=1.$
$\operatorname{rk}(\bigwedge^{2}E)=$ $\,\frac{r(r-1)}{2}$
$\operatorname{ch}_{1}(\bigwedge^{2}E)=$ $\,(r-1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\bigwedge^{2}E)=$ $\,(r-2)\operatorname{ch}_{2}(E)+\frac{1}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\bigwedge^{2}E)=$ $\,(r-4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{4}(\bigwedge^{2}E)=$ $\,(r-8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{1}{2}\operatorname{ch}_{2}(E)^{2}.$
@@
In this case we have three cases, namely $\alpha=(3),(2,1),(1,1,1)$.
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}+2^{a-1}+2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=1,$
@@
$\operatorname{rk}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+2}{3}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+2}{2}\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+3}{2}\operatorname{ch}_{2}(E)+\frac{r+2}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{(r+3)(r+6)}{2}\operatorname{ch}_{3}(E)+(r+3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{1}{6}\operatorname{ch}_{1}(E)^{3}$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{r^{2}+17r+54}{2}\operatorname{ch}_{4}(E)+(r+5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{r+4}{2}\operatorname{ch}_{2}(E)^{2}+\frac{1}{2}\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{(1,1)}-3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=2\mathsf{\mathbf{s}}_{1}\qquad\mbox{and}\qquad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=2,$
@@
$\operatorname{rk}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,\frac{r(r^{2}-1)}{3}$
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-1)\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-3)\operatorname{ch}_{2}(E)+r\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-9)\operatorname{ch}_{3}(E)+2r\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{1}(E)^{3}}{3}$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-27)\operatorname{ch}_{4}(E)+2r\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+r\operatorname{ch}_{2}(E)^{2}+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1,1)})=\mathsf{\mathbf{s}}_{(1,1)}-\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}-2^{a-1}-2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{c}\mathsf{D}_{b}\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=1,$
from which we deduce the first values of $\operatorname{ch}_{n}(\bigwedge^{3}E)$:
$\operatorname{rk}(\bigwedge^{3}E)=$ $\,\binom{r}{3}$
$\operatorname{ch}_{1}(\bigwedge^{3}E)=$ $\,\binom{r-1}{2}\operatorname{ch}_{1}(E)$
$\operatorname{ch}_{2}(\bigwedge^{3}E)=$ $\,\binom{r-2}{2}\operatorname{ch}_{2}(E)+\frac{r-2}{2}\operatorname{ch}_{1}(E)^{2}$
$\operatorname{ch}_{3}(\bigwedge^{3}E)=$ $\,\frac{(r-3)(r-6)}{2}\operatorname{ch}_{3}(E)+(r-3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{1}{6}\operatorname{ch}_{1}(E)^{3}$
$\operatorname{ch}_{4}(\bigwedge^{3}E)=$ $\,\frac{r^{2}-17r+54}{2}\operatorname{ch}_{4}(E)+(r-5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{r-4}{2}\operatorname{ch}_{2}(E)^{2}+\frac{1}{2}\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}.$
3.1. Chern character of symmetric and exterior powers
It will be useful to introduce the following notations. If $\nu=(1^{\nu_{1}},\dots,\ell^{\nu_{\ell}})$ is a partition, then we denote by

### v2

@@ 3.1. Chern character
$\mathsf{C}^{E}\colon\Lambda=\mathbb{Q}[\mathsf{\mathbf{p}}_{1},\mathsf{\mathbf{p}}_{2},\dots]\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E).$ (3.1)
Let us check that $\mathsf{C}^{E}$ is a characteristic class as from Definition 1.1.
Lemma 3.1.
Set $\mathsf{q}=\{k^{n}\}_{n\geq 1,k\geq 1}$. Then $\mathsf{C}^{E}$ is a $\operatorname{H}^{2\ast}(X,\mathbb{Q})$-valued $\mathsf{q}$-characteristic class.
Proof.
The first three points of Definition 1.1 are straightforward; notice that in this case $\mathsf{C}^{E}$ is fully additive, not just same-degree additive. Finally, the equality $\mathsf{C}_{n}^{E}(\mathsf{\mathbf{p}}_{k})=k^{n}\mathsf{C}_{n}^{E}(\mathsf{\mathbf{p}}_{1})$ can be checked by hand, or one can look at [DFO25a, Example 1.5]. ∎
Remark 3.2.
The set $\mathsf{q}=\{k^{n}\}_{n\geq 1,k\geq 1}$ is the same used to define the derivation $\mathsf{D}_{n}$ in § 2.3, see Remark 2.1.
Remark 3.3.
The proof of (4) in Definition 1.1 makes implicit use of the splitting principle: this is what we need in order to apply [DFO25a, Example 1.5]. We stress the fact that this is the only point where the splitting principle is used.
Recall that, by the hook length dimension formula, one has
@@
so that, for instance,
$\operatorname{rk}(\mathsf{\mathbf{S}}^{(2,1)}E)=\frac{r}{3}\cdot\frac{r+1}{1}\cdot\frac{r-1}{1}=2\binom{r+1}{3}\quad\mbox{ and }\quad\operatorname{rk}(\mathsf{\mathbf{S}}^{(3,2)}E)=\frac{(r+2)(r+1)r^{2}(r-1)}{24}.$
@@
The claim follows now by noticing that $\sum_{k}\mathsf{\mathbf{p}}_{k}{\partial}/{\partial\mathsf{\mathbf{p}}_{k}}$ acts on homogeneous elements via multiplication by their degree. ∎
We wish to point out the effectiveness of Theorem 3.4 by explicitly computing some examples.
@@
Let us compute $\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E)$ when $|\alpha|=2$. We have only two cases, namely the symmetric power $\mathsf{\mathbf{S}}^{2}E$ and the exterior power $\mathop{\bigwedge\nolimits^{\!2}}E$, corresponding to $\alpha=(1,1)$.
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2})=\mathsf{\mathbf{s}}_{1}+2^{a-1}\implies\operatorname{rk}(\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2}))=r+2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{a}\mathsf{D}_{b}(\mathsf{\mathbf{s}}_{2})=1.$
@@
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+1)\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\operatorname{ch}_{2}(E)+\frac{\operatorname{ch}_{1}(E)^{2}}{2},$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{2}(E)^{2}}{2},\dots$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1)})=\mathsf{\mathbf{s}}_{1}-2^{a-1}\implies\operatorname{rk}(\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{2}))=r-2^{a-1}\qquad\mbox{ and }\qquad\mathsf{D}_{a}\mathsf{D}_{b}(\mathsf{\mathbf{s}}_{(1,1)})=1.$
Then the first values of $\operatorname{ch}_{n}(\mathop{\bigwedge\nolimits^{\!2}}E)$ are:
$\operatorname{ch}_{1}(\mathop{\bigwedge\nolimits^{\!2}}E)=$ $\,(r-1)\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{2}(\mathop{\bigwedge\nolimits^{\!2}}E)=$ $\,(r-2)\operatorname{ch}_{2}(E)+\frac{\operatorname{ch}_{1}(E)^{2}}{2},$
$\operatorname{ch}_{3}(\mathop{\bigwedge\nolimits^{\!2}}E)=$ $\,(r-4)\operatorname{ch}_{3}(E)+\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{4}(\mathop{\bigwedge\nolimits^{\!2}}E)=$ $\,(r-8)\operatorname{ch}_{4}(E)+\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{2}(E)^{2}}{2},\dots$
@@
In this case we have three cases, namely $\alpha=(3),(2,1)$ and $(1,1,1)$.
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{a}\mathsf{D}_{b}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}+2^{a-1}+2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{a}\mathsf{D}_{b}\mathsf{D}_{c}(\mathsf{\mathbf{s}}_{3})=1,$
@@
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+2}{2}\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\binom{r+3}{2}\operatorname{ch}_{2}(E)+(r+2)\frac{\operatorname{ch}_{1}(E)^{2}}{2},$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{(r+3)(r+6)}{2}\operatorname{ch}_{3}(E)+(r+3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{1}(E)^{3}}{6},$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{3}E)=$ $\,\frac{r^{2}+17r+54}{2}\operatorname{ch}_{4}(E)+(r+5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+(r+4)\frac{\operatorname{ch}_{2}(E)^{2}}{2}+\frac{\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}}{2}.$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(2,1)})=\mathsf{\mathbf{s}}_{2}+\mathsf{\mathbf{s}}_{(1,1)}-3^{a-1},\quad\mathsf{D}_{a}\mathsf{D}_{b}(\mathsf{\mathbf{s}}_{(2,1)})=2\mathsf{\mathbf{s}}_{1}\qquad\mbox{and}\qquad\mathsf{D}_{a}\mathsf{D}_{b}\mathsf{D}_{c}(\mathsf{\mathbf{s}}_{(2,1)})=2,$
@@
$\operatorname{ch}_{1}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-1)\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{2}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-3)\operatorname{ch}_{2}(E)+2r\frac{\operatorname{ch}_{1}(E)^{2}}{2},$
$\operatorname{ch}_{3}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-9)\operatorname{ch}_{3}(E)+2r\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+2\frac{\operatorname{ch}_{1}(E)^{3}}{6},$
$\operatorname{ch}_{4}(\mathsf{\mathbf{S}}^{(2,1)}E)=$ $\,(r^{2}-27)\operatorname{ch}_{4}(E)+2r\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+2r\frac{\operatorname{ch}_{2}(E)^{2}}{2}+2\frac{\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}}{2},\dots$
@@
$\mathsf{D}_{a}(\mathsf{\mathbf{s}}_{(1,1,1)})=\mathsf{\mathbf{s}}_{(1,1)}-\mathsf{\mathbf{s}}_{1}2^{a-1}+3^{a-1},\quad\mathsf{D}_{a}\mathsf{D}_{b}(\mathsf{\mathbf{s}}_{3})=\mathsf{\mathbf{s}}_{1}-2^{a-1}-2^{b-1}\quad\mbox{and}\quad\mathsf{D}_{a}\mathsf{D}_{b}\mathsf{D}_{c}(\mathsf{\mathbf{s}}_{3})=1,$
from which we deduce the first values of $\operatorname{ch}_{n}(\mathop{\bigwedge\nolimits^{\!3}}E)$:
$\operatorname{ch}_{1}(\mathop{\bigwedge\nolimits^{\!3}}E)=$ $\,\binom{r-1}{2}\operatorname{ch}_{1}(E),$
$\operatorname{ch}_{2}(\mathop{\bigwedge\nolimits^{\!3}}E)=$ $\,\binom{r-2}{2}\operatorname{ch}_{2}(E)+(r-2)\frac{\operatorname{ch}_{1}(E)^{2}}{2},$
$\operatorname{ch}_{3}(\mathop{\bigwedge\nolimits^{\!3}}E)=$ $\,\frac{(r-3)(r-6)}{2}\operatorname{ch}_{3}(E)+(r-3)\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)+\frac{\operatorname{ch}_{1}(E)^{3}}{6},$
$\operatorname{ch}_{4}(\mathop{\bigwedge\nolimits^{\!3}}E)=$ $\,\frac{r^{2}-17r+54}{2}\operatorname{ch}_{4}(E)+(r-5)\operatorname{ch}_{3}(E)\operatorname{ch}_{1}(E)+(r-4)\frac{\operatorname{ch}_{2}(E)^{2}}{2}+\frac{\operatorname{ch}_{2}(E)\operatorname{ch}_{1}(E)^{2}}{2},\dots$
3.2. Twisted Chern character
In [Mar20], Markman introduces the twisted Chern character of a vector bundle $F$ as
This class should be thought of as the Chern character of the projective bundle $\mathbb{P}E$, which motivates the terminology. For bookkeeping, let us express the first three values of $k(F)$ in terms of Chern classes:
and
Notice that $2\operatorname{rk}(F)\,k_{2}(F)=\Delta(F)$ is the discriminant of $F$.
Define
$\mathsf{C}^{\mathbb{P}E}\colon\Lambda\to\operatorname{H}^{2\ast}(X,\mathbb{Q}),\qquad\mathsf{\mathbf{s}}_{\alpha}\mapsto k(\mathsf{\mathbf{S}}^{\alpha}E)=\exp\left(-\frac{|\alpha|}{r}\operatorname{ch}_{1}(E)\right)\operatorname{ch}(\mathsf{\mathbf{S}}^{\alpha}E),$
where the equality follows from Corollary 3.5.
Lemma 3.8.
Set $\mathsf{q}=\{k^{n}\}_{n\geq 1,k\geq 1}$. Then $\mathsf{C}^{\mathbb{P}E}$ is a $\operatorname{H}^{2\ast}(X,\mathbb{Q})$-valued $\mathsf{q}$-characteristic class.
Proof.
The first two properties of Definition 1.1 are straightforward. Moreover, $\mathsf{C}^{\mathbb{P}E}$ is same-degree additive, since the degree $d$ homogeneous component of $\Lambda$ is generated by Schur functions $\mathsf{\mathbf{s}}_{\alpha}$ with $|\alpha|=d$.
Finally, let us notice that $\mathsf{C}^{\mathbb{P}E}_{1}(\mathsf{\mathbf{p}}_{k})=0=\mathsf{C}^{\mathbb{P}E}(\mathsf{\mathbf{p}}_{1})$, while $\mathsf{C}^{\mathbb{P}E}_{n}(\mathsf{\mathbf{p}}_{k})=k^{n}\mathsf{C}^{\mathbb{P}E}_{n}(\mathsf{\mathbf{p}}_{1})$, as it can be computed by hand using the same equality for $\mathsf{C}^{E}$. ∎
As a direct application of Corollary 1.4, we get the following formula for the class $k$ of Schur bundles.
Theorem 3.9.
With notations as above,
Example 3.10.
so that, for instance,
$k_{2}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+2)\,k_{2}(E)\,$,
$k_{3}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+4)\,k_{3}(E)\,$,
$k_{4}(\mathsf{\mathbf{S}}^{2}E)=$ $\,(r+8)\,k_{4}(E)+\frac{1}{2}k_{2}^{2}(E)\,.$
Remark 3.11.
from which it follows, for example, that
In [DFO25a] we proved that the classes $\mathsf{L}_{2}(F)$ and $\mathsf{L}_{3}(F)$ are eigen-classes with respect to the induced action of Schur functors. Moreover, we also remarked that this claim does not extend to higher-degree classes and speculated that similar formulas should exist for the projection of the classes $\mathsf{L}_{n}(\mathsf{\mathbf{S}}^{\alpha}E)$ onto the $\mathsf{L}_{n}(E)$-component.
Remark 3.12.
Since $\mathsf{C}^{\mathbb{P}E}_{1}(\mathsf{\mathbf{p}}_{k})=0=\mathsf{C}^{\mathbb{P}E}(\mathsf{\mathbf{p}}_{1})$, the numbers $q_{1,k}$ can be arbitrarily chosen. Any special choice of such numbers has the effect of modifying the single derivation $\mathsf{D}_{1}$, which is however only involved in those terms $k^{(\nu)}$ containing $k_{1}(E)=0$ that do, in any case, vanish altogether.
4. Chern characters of symmetric and exterior powers
In this section we explicitly compute the (twisted) Chern character of the symmetric and exterior powers of $E$, recovering precisely the formulas obtained by Manivel in [Man94] in the symmetric case and finding expressions equivalent to those by Svrtan in [Svr93], which use Stirling numbers of the second kind instead of Eulerian numbers.
Let us start by setting notations. If $\nu=(1^{\nu_{1}},\dots,\ell^{\nu_{\ell}})$ is a partition, then we denote by